OpenStudy (anonymous):

Two more! 1) What is the solution set of |2x + 4| < 8? AND 2) What is the solution set of |4x – 3| – 2 > 3?

hero (hero):

1. -8 < 2x+4 < 8

OpenStudy (anonymous):

1) \[-6 < x < 2\] 2) \[2 < x , x < -0.5\]

OpenStudy (anonymous):

Thank you! :D

OpenStudy (ybarrap):

|4x -3| - 2 > 3 |4x -3| > 3 + 2 = 5 4x -3 > 5 and 4x -3 < -5 4x > 5 + 3 =8 and 4x < -5 +3 = -2 x > 8/4 = 2 and x < -2/4 = -1/2 Ans. x > 2 and x < -1/2